OpenStudy (anonymous):

Using complete sentences, explain what the discriminant is and what it tells you about the solutions of a quadratic equation. Provide a unique example to back up your explanation.

jhonyy9 (jhonyy9):

when the discriminant is greater than zero D>0 there are two real solutions differents - when D=0 there are two solutions equales - when D<0 there are two solutions complex

jhonyy9 (jhonyy9):

1. \[x ^{2} +4x +3 =0\]

jhonyy9 (jhonyy9):

D=16-12=4 x_1,_2=(-4+/-sqrt(4))/2=(-4+/-2)/2= -3 and -1

jhonyy9 (jhonyy9):

2.

jhonyy9 (jhonyy9):

\[x ^{2}+2x+1=0\]

jhonyy9 (jhonyy9):

D=4-4=0 x_1,_2=(-2+/- 0)/2= -2/2 =-1

jhonyy9 (jhonyy9):

3.

jhonyy9 (jhonyy9):

\[x ^{2}-x+1=0\]

jhonyy9 (jhonyy9):

D=1-4= -3 x_1,_2=(1+/- sqrt(-3))/2 = (1+/- (i)sqrt(3))/2